• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 16:28:04 on June 5,2005

    Entry number 145157

    keyword=day shift summary

    started day shift with magnets dropping to 1.850 GeV/c.

    LeRose has come in and noted that people have (incorrectly) been using set currents on septa, not readback. We have decided to continue this for now, since most data was taken like this. It affects the y-position mostly.

    Also messing with run start and stop scripts to add septum and SLI information.

    Run 4274: P=1.887 (+2%), thin carbon

    change to 1.850

    Run 4275: P=1.850 (0%), thin carbon
    Run 4276: junk
    Run 4277: P=1.850 (0%), tantalum

    change to 1.813

    Run 4278: P=1.813 (-2%), tantalum, lots of cosmics (beam gone awhile)
    Run 4279: P=1.813 (-2%), tantalum, repeat
    Run 4280: P=1.813 (-2%), thin carbon

    change to 1.776

    run 4283 tantalum for Rob

    change to 1.906 (+3.027%)

    run 4284 tantalum

    move target on thin C

    run 4286 on thin Carbon

    change to 1.867 (+0.919%)

    run 4287 on thin Carbon, lost beam, end-run failed

    move target on tantalum

    run 4288 on tantalum

    change to 1.832 (-0.98%)

    run 4289 on tantalum